OpenStudy (anonymous):

Select the two values of x that are roots of the given polynomial below. x^2=3x-5

OpenStudy (mathmale):

Welcome to OpenStudy, pashes! Although you could just try the given x values in the list of possible answers, why not get practice in actually solving quadratics? If x^2=3x-5, then \[x^2- 3x + 5 = 0\]

OpenStudy (mathmale):

Unfortunately, this doesn't have nice integer roots; that means we'll need to use the Quadratic formula to find the two roots. Have you used the Q. F. before? Note that in \[x^2- 3x + 5 = 0\] a=1, b=-3, and c=5.